OpenStudy (anonymous):

Solve, round your answer to the nearest 2 decimal places 2x4 = 162

OpenStudy (unklerhaukus):

Do you mean\[2x^4=16^2\]?

OpenStudy (anonymous):

yea but its =162

OpenStudy (unklerhaukus):

ok \[2x^4=162\]

OpenStudy (unklerhaukus):

First step to to get rid of the two, divide both sides of the equation by 2

OpenStudy (unklerhaukus):

\[2x^4/2=162/2\] what does this simplify to become?

OpenStudy (anonymous):

81 is what i got i divide 162

OpenStudy (unklerhaukus):

\[2x^4/2=81\]

OpenStudy (anonymous):

thats the answer

OpenStudy (unklerhaukus):

not yet, the answer will be \[x= \color{brown}{\text{something}}\] we have to simplify the left hand side, by cancelling the twos and then taking the forth root

OpenStudy (anonymous):

x=9

OpenStudy (unklerhaukus):

\[x^4=81\\x=\sqrt[4]{81}\]

OpenStudy (anonymous):

so x=81

OpenStudy (unklerhaukus):

can you express 81 as a square number?

OpenStudy (anonymous):

idk

OpenStudy (unklerhaukus):

\[a^2 = 81\] what is a?

OpenStudy (unklerhaukus):

(check your times tables)

OpenStudy (anonymous):

im new to this

OpenStudy (unklerhaukus):

\[a^2 =a\times a=81\]

OpenStudy (unklerhaukus):

Are you still here @morrisd7 ?

OpenStudy (anonymous):

yea i got 3 min left on this test help plz